OpenStudy (anonymous):

\[A= (a ^{-2}*a ^{4})^{-1}*(a ^{-2})^{-3}*(a ^{7})^{-1}\]

OpenStudy (anonymous):

\[A=(a^{2})^{-1}\times(a)^{6}\times(a)^{-7}\] \[A=(a)^{-2*6*−7}\] A=a^84

OpenStudy (anonymous):

second line: wrong: A=a^(-2+6-7)=a^(-3) :))

OpenStudy (anonymous):

is that anwser simplifed?